OpenStudy (anonymous):

Determine the sum of the measures of the interior angles of the specified polygon. An octagon A. 720° B. 1080° C. 900° D. 1440°

OpenStudy (campbell_st):

use the formula Angle Sum = (n -2) x 180 n = number of sides

OpenStudy (anonymous):

So when you plug it in it would = (8-2) x 180 = 1080